OpenStudy (anonymous):

The number of pets owned by eight classmates is shown below: 1, 3, 2, 2, 1, 3, 3, 1 Part A: What is the mean of the data? Show your work. (4 points) Part B: Use your answer from Part A to calculate the mean absolute deviation for the data. Show your work. (6 points)

OpenStudy (anonymous):

I just need help with b

OpenStudy (anonymous):

I got 2 for the absolute deviation is that correct?

OpenStudy (anonymous):

what up

OpenStudy (anonymous):

it is 2

OpenStudy (anonymous):

so you was correct

OpenStudy (anonymous):

can you show me how you know?

OpenStudy (welshfella):

what did you get for the mean?

OpenStudy (anonymous):

first you add all of the numbers which was 16 and then divide by how many people there was which was 8 so 16 divided by 8 is 2

OpenStudy (welshfella):

thats the mean 2

OpenStudy (anonymous):

for the part a i got 2

OpenStudy (welshfella):

to get the mean absolute deviation you add up the absolute differences of each value from the mean . Add these values then divide by 8

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

1 2-1 = 1 3 3 -2 = 1 2 2 -2 = 0 2 2 - 2 = 0 and so on

OpenStudy (anonymous):

i did this to find the absolute deviation 1-2+3-2+2-2+2-2+1-2+3-2+3-2+1-2

OpenStudy (anonymous):

so all that above 8

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

so 6/8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (welshfella):

Hmm to be honest i'm not sure I'll have to check

OpenStudy (anonymous):

same im confused im going to go over it again and you tell me how you it so we can compare

OpenStudy (anonymous):

*do

OpenStudy (welshfella):

right - you take ABSOLUTE values so for instance first values is 1: 1 - 2 = -1 The absolute value of -1 is 1. so there are no negatives

OpenStudy (welshfella):

so 6/8 is not correct

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

i was doing something else what

OpenStudy (welshfella):

oh hold on it does work out to 6/8 it is 1+ 1 +0+ 0 + 1+ 1+1 + 1 = 6 so 6/8 or 3/4

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

what you wrote 1-2+3-2+2-2+2-2+1-2+3-2+3-2+1-2 = 0